OpenStudy (anonymous):

What is dy/dx of y=ln(5-x^2)??

OpenStudy (anonymous):

dy/dx = -2x/(5-x^2)

OpenStudy (anonymous):

The derivative of a logarithmic function y = ln(u) where u is a function: dy/dx = (1/u)(du/dx)

OpenStudy (anonymous):

u=5-x^2 y=log(u) the chain rule says: \[dy/dx = (dy/du)(du/dx)\] dy/du=-2x du/dx=1/u finally. dy/dx = (-2x)(1/u) replacing u dy/dx = (-2x)(1/(5-x^2)) dy/dx = 2x/(x^2 - 5)

OpenStudy (anonymous):

any question?

OpenStudy (anonymous):

nope thats great thanks!

OpenStudy (anonymous):

np